OpenStudy (anonymous):

Fun Land theme park has a total of 80 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? 80 60 30 20

OpenStudy (anonymous):

Females—x: x + 2x + 20 = 110 3x = 90 x = 30 Males: = 2(30) + 20 = 60 + 20 = 80 Answer (1): 80 males Checking: 80 - 2(110 - 80) = 20 80 - 2(30) = 20 80 - 60 = 20 20 = 20 ----------- No. of men—x: 2x - 3(100 - x) = 20 2x - 300 + 3x = 20 5x = 320 x = 64 Answer (2): 64 men Checking: 2(64) - 3(100 - 64) = 20 128 - 3(36) = 20 128 - 108 = 20 20 = 20

OpenStudy (whpalmer4):

@Bri_Love were you doing a different problem as an example?

OpenStudy (whpalmer4):

Problem statement says that there are a total of 80 employees, not 110.

OpenStudy (anonymous):

Nope that should be the answer.

OpenStudy (anonymous):

Its 20

OpenStudy (whpalmer4):

No. M = number of males F = number of females M+F = 80 M = 2*F + 20 2*F + 20 + F = 80 3*F = 60 F = 20 M = 80-F = 80 - 20 = 60

OpenStudy (anonymous):

ohhhh yeah i messed up @whpalmer4 your right is 60

OpenStudy (whpalmer4):

Always good to reread the problem with answer in hand to make sure it fits!

OpenStudy (whpalmer4):

I like the male/female breakdown in your answer better, though :-)